OpenStudy (anonymous):

solve for x e^(3-4x) = 5

OpenStudy (anonymous):

x=l(n(5)-3)/(-4)

OpenStudy (anonymous):

x=(ln(5)-3)/(-4)= 0.3476...

OpenStudy (anonymous):

i got lne^(3-4x)=ln5

OpenStudy (saifoo.khan):

right.

OpenStudy (anonymous):

so shouldnt it be 3-5x=ln5

OpenStudy (saifoo.khan):

3-4x = ln5

OpenStudy (anonymous):

what im not sure is to do next

OpenStudy (saifoo.khan):

3-4x = 1.609

OpenStudy (saifoo.khan):

Solve for x.

OpenStudy (anonymous):

My teacher is looking for exact values

OpenStudy (anonymous):

-ln5+3 / 4 = x

OpenStudy (anonymous):

would that be right?

OpenStudy (saifoo.khan):

i guess so...

OpenStudy (anonymous):

so ln would be the more appropriate option than using log